# Accurate and thermodynamically consistent hydrogen equation of state for planetary modeling with flow matching

### v1

@@ Acknowledgements.
- Satorras et al. [2021] V. G. Satorras, E. Hoogeboom, and M. Welling, E(n) equivariant graph neural networks, in Proceedings of the 38th International Conference on Machine Learning, Proceedings of Machine Learning Research, Vol. 139, edited by M. Meila and T. Zhang (PMLR, 2021) pp. 9323–9332.
@@
- Taddei et al. [2015] M. Taddei, M. Ruggeri, S. Moroni, and M. Holzmann, Iterative backflow renormalization procedure for many-body ground-state wave functions of strongly interacting normal fermi liquids, Phys. Rev. B 91, 115106 (2015).
@@
- Pfau et al. [2020] D. Pfau, J. S. Spencer, A. G. D. G. Matthews, and W. M. C. Foulkes, Ab initio solution of the many-electron schrödinger equation with deep neural networks, Phys. Rev. Res. 2, 033429 (2020).
@@
- Pescia et al. [2022] G. Pescia, J. Han, A. Lovato, J. Lu, and G. Carleo, Neural-network quantum states for periodic systems in continuous space, Phys. Rev. Res. 4, 023138 (2022).
@@
- Wilson et al. [2023] M. Wilson, S. Moroni, M. Holzmann, N. Gao, F. Wudarski, T. Vegge, and A. Bhowmik, Neural network ansatz for periodic wave functions and the homogeneous electron gas, Phys. Rev. B 107, 235139 (2023).
@@
- Cassella et al. [2023] G. Cassella, H. Sutterud, S. Azadi, N. D. Drummond, D. Pfau, J. S. Spencer, and W. M. C. Foulkes, Discovering quantum phase transitions with fermionic neural networks, Phys. Rev. Lett. 130, 036401 (2023).
@@
- Pescia et al. [2024] G. Pescia, J. Nys, J. Kim, A. Lovato, and G. Carleo, Message-passing neural quantum states for the homogeneous electron gas, Phys. Rev. B 110, 035108 (2024).
@@
- Giannozzi et al. [2009] P. Giannozzi, S. Baroni, N. Bonini, M. Calandra, R. Car, C. Cavazzoni, D. Ceresoli, G. L. Chiarotti, M. Cococcioni, I. Dabo, A. Dal Corso, S. de Gironcoli, S. Fabris, G. Fratesi, R. Gebauer, U. Gerstmann, C. Gougoussis, A. Kokalj, M. Lazzeri, L. Martin-Samos, N. Marzari, F. Mauri, R. Mazzarello, S. Paolini, A. Pasquarello, L. Paulatto, C. Sbraccia, S. Scandolo, G. Sclauzero, A. P. Seitsonen, A. Smogunov, P. Umari, and R. M. Wentzcovitch, Quantum espresso: a modular and open-source software project for quantum simulations of materials, Journal of Physics: Condensed Matter 21, 395502 (19pp) (2009).
- Blöchl [1994] P. E. Blöchl, Projector augmented-wave method, Phys. Rev. B 50, 17953 (1994).
@@
- Monkhorst and Pack [1976] H. J. Monkhorst and J. D. Pack, Special points for brillouin-zone integrations, Phys. Rev. B 13, 5188 (1976).
@@
- Bussi et al. [2007] G. Bussi, D. Donadio, and M. Parrinello, Canonical sampling through velocity rescaling, The Journal of Chemical Physics 126, 014101 (2007).
@@
- Chen and Lipman [2024] R. T. Q. Chen and Y. Lipman, Flow matching on general geometries, in The Twelfth International Conference on Learning Representations (2024).
@@
- Klein et al. [2023] L. Klein, A. Krämer, and F. Noe, Equivariant flow matching, in Thirty-seventh Conference on Neural Information Processing Systems (2023).
@@
- Kuhn [1955] H. W. Kuhn, The hungarian method for the assignment problem, Naval Research Logistics Quarterly 2, 83 (1955).
@@
- Militzer [2013] B. Militzer, Equation of state calculations of hydrogen-helium mixtures in solar and extrasolar giant planets, Phys. Rev. B 87, 014202 (2013).
@@
- Becker et al. [2013] A. Becker, N. Nettelmann, B. Holst, and R. Redmer, Isentropic compression of hydrogen: Probing conditions deep in planetary interiors, Phys. Rev. B 88, 045122 (2013).
@@
- Morales et al. [2013] M. A. Morales, J. M. McMahon, C. Pierleoni, and D. M. Ceperley, Nuclear quantum effects and nonlocal exchange-correlation functionals applied to liquid hydrogen at high pressure, Phys. Rev. Lett. 110, 065702 (2013).
@@
- Chabrier and Potekhin [1998] G. Chabrier and A. Y. Potekhin, Equation of state of fully ionized electron-ion plasmas, Phys. Rev. E 58, 4941 (1998).
@@
- Militzer and Ceperley [2000] B. Militzer and D. M. Ceperley, Path integral monte carlo calculation of the deuterium hugoniot, Phys. Rev. Lett. 85, 1890 (2000).
@@
- Militzer and Ceperley [2001] B. Militzer and D. M. Ceperley, Path integral monte carlo simulation of the low-density hydrogen plasma, Phys. Rev. E 63, 066404 (2001).
@@
In this work, we use the flow matching method to benchmark the calculation of free energy difference (scaled by temperature) between the two states, , against the conventional thermodynamic integration (TI) approach. Below we will discuss various aspect of the new method, including the underlying theoretical framework, training objectives, implementation details, and some illustrative examples of the results.
@@ Appendix A Targeted free energy perturbation
In contrast to TI, flow matching is built on the framework of targeted free energy perturbation (TFEP) [28, 29], as already mentioned in the main text. Compared to the original free energy perturbation based on importance sampling [58], TFEP further leverages an invertible map in the configuration space to significantly increase state overlap and thereby reduces the variance of relevant estimators.
@@
Note that due to limited number of samples in practice, the two estimators in Eq. (S4) are actually biased, which can be readily demonstrated using Jensen’s inequality [29]. Specifically, the estimator involving the forward (reverse) work turns out to be an upper (lower) bound of the true free energy difference value. The extent of bias will, of course, depend on the quality of the map . In this work, we found these two bounds are generally quite tight; see the sections below for some illustrative examples. As a result, we choose to simply perform an arithmetic mean of the two bounds to generate the final data, although some more sophisticated and potentially superior approaches may also be considered [59]. Another observation is that the estimation of absolute entropy appears to be more straightforward within the TFEP framework compared to TI: once one of the two states is chosen to have tractable entropy, the practical implementation remains almost unchanged.
@@ Appendix B Flow matching
The recent progress in the field of deep generative modeling has largely changed the situation. Specifically, the desired invertible map can be efficiently represented by a class of generative models called normalizing flow [30], making TFEP reemerge as a promising tool for free energy calculation. In previous attempts [31, 32, 33, 34], the training objectives have been exclusively based on variants of the Kullback-Leibler (KL) divergence, which provide direct access to the free energy difference as a variational upper or lower bound. This variational training scheme has been successfully applied to a broad range of physical problems, including lattice models [35, 36], atomic solids [32, 33], quantum dots [37], uniform electron gases [38], and also dense hydrogen [39]. However, it still has some drawbacks for the present purpose. First, the training process typically requires a large number of energy calculations of the underlying system, which would be time-consuming for ab initio methods like DFT. As a result, many works have used empirical [31, 32, 33] or machine-learned potentials [34], which would inevitably introduce new source of errors. Second, symmetry constraints often play an essential role in the efficient training of flow model, especially for systems like pure hydrogen. Although several neural network designs have been proposed to meet various symmetry requirements [31, 32], they are still not satisfactory enough regarding a proper balance between the expressive power and computational cost. The continuous normalizing flow (CNF) [60, 61] is arguably the most convenient and flexible candidate now to incorporate symmetries, but its variational training requires expensive numerical integrations of ordinary differential equation (ODE) and scales unfavorably for large system sizes.
The recently developed flow matching method [23, 24, 25] has largely overcome these difficulties. More specifically, one can smoothly connect the two end probability distributions by introducing an interpolant , where is a continuous “time” parameter, such as
@@
Notice this expression requires only samples from the two end distributions; there are not any on-the-fly evaluation of configurational energy or ODE integration during the training stage. On the other hand, these types of costly computations are indeed present when estimating the free energy difference after the training is finished, as shown in Eqs. (S4) and (S5). In particular, the transformed coordinates and corresponding log-Jacobian determinant should be evaluated by jointly integrating the following set of ODEs [60, 61]:
@@
As mentioned above, compared to previous network designs for discrete flow model [31, 32], the incorporation of various symmetries can also be more convenient for continuous normalizing flow. Specifically, the invariance of probability distribution can simply be achieved by requiring the underlying velocity field to be equivariant under the desired symmetry operation :
To do this in practice, we can leverage many recent advances in various areas, such as natural language processing [62], molecular simulation [63, 31], graph neural networks [64], and quantum many body computations [65, 66, 38, 67, 39, 68, 69, 70].
@@ Appendix C Implementation details
The MD generation of coordinate samples used for training Eq. (S7) and the energy computation appearing in the free energy estimators Eq. (S4) are both performed in the DFT level using the QUANTUM ESPRESSO code [71]. We use the exchange-correlation functional of Perdew, Burke and Ernzerhof (PBE) [15] and the pseudopotential of projector-augmented wave [72] type. We choose the system size , and the -space integration is performed on a Monkhorst-Pack grid [73]. The energy cutoff for the plane wave basis set and charge densities are and Ry, respectively. These parameters have been checked to yield well converged energetics. For MD simulations, the ion temperature is controlled using stochastic-velocity rescaling [74] and the time step is set to 12 a.u..
The network we use for the velocity field is adapted from FermiNet [66], which is permutation and translation equivariant by construction. We also modify the pair distance features to comply with the periodic nature of the simulation box [38, 67, 68, 69]. We choose a simple linear interpolating function
with the caveat in mind that we are working on a periodic box. This setting turns out to be equivalent to the Riemannian flow matching on a torus as proposed in Ref. [75]. To allow for more efficient use of the training data, we choose to further incorporate permutation symmetry explicitly into the flow matching objective Eq. (S7) itself [76]. In practice, this can be achieved by finding the permutation with minimal distance for each pair of samples using the Hungarian algorithm [77].
@@ Appendix D Illustrative examples of the results
After the training is complete, one can further use Eq. (S4) to estimate the upper and lower bound of free energy difference between the two states, as shown by the red and black vertical line in Fig. S2, respectively. The distance between these two bounds is roughly within their own error bars, which implies our estimation have been tight enough. Figure S2 also shows histograms for the distribution of forward and reverse work , as defined in Eq. (S5). The significant overlap between these two work distributions indicates the high quality of the trained velocity field in effectively connecting the two end states being considered. Furthermore, one can clearly see the estimated free energy difference values coincide well with the intersection point of the two work distributions, where the probability densities are equal. This is a direct consequence of the fluctuation theorem as derived in Ref. [29].
We also use the same workflow outlined above to calculate the absolute free energy and entropy of the reference state at and , as mentioned in the main text. In practice, we choose the “tractable state” to correspond to a simple uniform distribution. The resulting radial distribution functions and free energy bounds are shown in Figures S3 and S4, respectively. Notice that in the present case, the two end states differ more substantially than those in Fig. S1, making the flow matching training of the velocity field more challenging. As a result, the forward and reverse work distributions also have less overlap, and the upper and lower bounds for the free energy difference are less tight compared to Fig. S2. Nonetheless, the absolute entropy of the reference state can still be estimated with a relative error of , which is comparable to the results reported by Morales et al. [19] using Hamiltonian thermodynamic integration and already satisfactory for the present purpose (see Fig. 1(a) in the main text).

### v2

@@ Acknowledgements.
- Satorras et al. (2021) Víctor Garcia Satorras, Emiel Hoogeboom, and Max Welling, “E(n) equivariant graph neural networks,” in Proceedings of the 38th International Conference on Machine Learning, Proceedings of Machine Learning Research, Vol. 139, edited by Marina Meila and Tong Zhang (PMLR, 2021) pp. 9323–9332.
@@
- Taddei et al. (2015) Michele Taddei, Michele Ruggeri, Saverio Moroni, and Markus Holzmann, “Iterative backflow renormalization procedure for many-body ground-state wave functions of strongly interacting normal fermi liquids,” Phys. Rev. B 91, 115106 (2015).
@@
- Pfau et al. (2020) David Pfau, James S. Spencer, Alexander G. D. G. Matthews, and W. M. C. Foulkes, “Ab initio solution of the many-electron schrödinger equation with deep neural networks,” Phys. Rev. Res. 2, 033429 (2020).
@@
- Pescia et al. (2022) Gabriel Pescia, Jiequn Han, Alessandro Lovato, Jianfeng Lu, and Giuseppe Carleo, “Neural-network quantum states for periodic systems in continuous space,” Phys. Rev. Res. 4, 023138 (2022).
@@
- Wilson et al. (2023) Max Wilson, Saverio Moroni, Markus Holzmann, Nicholas Gao, Filip Wudarski, Tejs Vegge, and Arghya Bhowmik, “Neural network ansatz for periodic wave functions and the homogeneous electron gas,” Phys. Rev. B 107, 235139 (2023).
@@
- Cassella et al. (2023) Gino Cassella, Halvard Sutterud, Sam Azadi, N. D. Drummond, David Pfau, James S. Spencer, and W. M. C. Foulkes, “Discovering quantum phase transitions with fermionic neural networks,” Phys. Rev. Lett. 130, 036401 (2023).
@@
- Pescia et al. (2024) Gabriel Pescia, Jannes Nys, Jane Kim, Alessandro Lovato, and Giuseppe Carleo, “Message-passing neural quantum states for the homogeneous electron gas,” Phys. Rev. B 110, 035108 (2024).
@@
- Giannozzi et al. (2009) Paolo Giannozzi, Stefano Baroni, Nicola Bonini, Matteo Calandra, Roberto Car, Carlo Cavazzoni, Davide Ceresoli, Guido L Chiarotti, Matteo Cococcioni, Ismaila Dabo, Andrea Dal Corso, Stefano de Gironcoli, Stefano Fabris, Guido Fratesi, Ralph Gebauer, Uwe Gerstmann, Christos Gougoussis, Anton Kokalj, Michele Lazzeri, Layla Martin-Samos, Nicola Marzari, Francesco Mauri, Riccardo Mazzarello, Stefano Paolini, Alfredo Pasquarello, Lorenzo Paulatto, Carlo Sbraccia, Sandro Scandolo, Gabriele Sclauzero, Ari P Seitsonen, Alexander Smogunov, Paolo Umari, and Renata M Wentzcovitch, “Quantum espresso: a modular and open-source software project for quantum simulations of materials,” Journal of Physics: Condensed Matter 21, 395502 (19pp) (2009).
- Blöchl (1994) P. E. Blöchl, “Projector augmented-wave method,” Phys. Rev. B 50, 17953–17979 (1994).
@@
- Monkhorst and Pack (1976) Hendrik J. Monkhorst and James D. Pack, “Special points for brillouin-zone integrations,” Phys. Rev. B 13, 5188–5192 (1976).
@@
- Bussi et al. (2007) Giovanni Bussi, Davide Donadio, and Michele Parrinello, “Canonical sampling through velocity rescaling,” The Journal of Chemical Physics 126, 014101 (2007).
@@
- Chen and Lipman (2024) Ricky T. Q. Chen and Yaron Lipman, “Flow matching on general geometries,” in The Twelfth International Conference on Learning Representations (2024).
@@
- Klein et al. (2023) Leon Klein, Andreas Krämer, and Frank Noe, “Equivariant flow matching,” in Thirty-seventh Conference on Neural Information Processing Systems (2023).
@@
- Kuhn (1955) H. W. Kuhn, “The hungarian method for the assignment problem,” Naval Research Logistics Quarterly 2, 83–97 (1955).
- Saumon and Chabrier (1991) Didier Saumon and Gilles Chabrier, “Fluid hydrogen at high density: Pressure dissociation,” Phys. Rev. A 44, 5122–5141 (1991).
@@
- Militzer (2013) Burkhard Militzer, “Equation of state calculations of hydrogen-helium mixtures in solar and extrasolar giant planets,” Phys. Rev. B 87, 014202 (2013).
@@
- Becker et al. (2013) Andreas Becker, Nadine Nettelmann, Bastian Holst, and Ronald Redmer, “Isentropic compression of hydrogen: Probing conditions deep in planetary interiors,” Phys. Rev. B 88, 045122 (2013).
@@
- Morales et al. (2013) Miguel A. Morales, Jeffrey M. McMahon, Carlo Pierleoni, and David M. Ceperley, “Nuclear quantum effects and nonlocal exchange-correlation functionals applied to liquid hydrogen at high pressure,” Phys. Rev. Lett. 110, 065702 (2013).
@@
- Chabrier and Potekhin (1998) Gilles Chabrier and Alexander Y. Potekhin, “Equation of state of fully ionized electron-ion plasmas,” Phys. Rev. E 58, 4941–4949 (1998).
@@
- Militzer and Ceperley (2000) B. Militzer and D. M. Ceperley, “Path integral monte carlo calculation of the deuterium hugoniot,” Phys. Rev. Lett. 85, 1890–1893 (2000).
@@
- Militzer and Ceperley (2001) B. Militzer and D. M. Ceperley, “Path integral monte carlo simulation of the low-density hydrogen plasma,” Phys. Rev. E 63, 066404 (2001).
@@
In this work, we use the flow matching method to benchmark the calculation of free energy difference (scaled by temperature) between the two states, , against the conventional thermodynamic integration (TI) approach. Below we will discuss various aspects of the new method, including the underlying theoretical framework, training objectives, implementation details, and some illustrative examples of the results.
@@ Appendix A Targeted free energy perturbation
In contrast to TI, flow matching is built on the framework of targeted free energy perturbation (TFEP) Jarzynski (2002); Hahn and Then (2009), as already mentioned in the main text. Compared to the original free energy perturbation based on importance sampling Zwanzig (1954), TFEP further leverages an invertible map in the configuration space to significantly increase state overlap and thereby reduces the variance of relevant estimators.
@@
Note that due to the limited number of samples in practice, the two estimators in Eq. (S4) are actually biased, which can be readily demonstrated using Jensen’s inequality Hahn and Then (2009). Specifically, the estimator involving the forward (reverse) work turns out to be an upper (lower) bound of the true free energy difference value. The extent of bias will, of course, depend on the quality of the map . In this work, we found these two bounds are generally quite tight; see the sections below for some illustrative examples. As a result, we choose to simply perform an arithmetic mean of the two bounds to generate the final free energy and entropy data, although some more sophisticated and potentially superior approaches may also be considered Bennett (1976). Another observation is that the estimation of absolute entropy appears to be more straightforward within the TFEP framework compared to TI: once one of the two states is chosen to have known entropy, the practical implementation remains almost unchanged.
@@ Appendix B Flow matching
The recent progress in the field of deep generative modeling has largely changed the situation. Specifically, the desired invertible map can be efficiently represented by a class of generative models called normalizing flow Papamakarios et al. (2021), making TFEP reemerge as a promising tool for free energy calculation. In previous attempts Wirnsberger et al. (2020, 2022); Ahmad and Cai (2022); Molina-Taborda et al. (2024), the training objectives have been exclusively based on variants of the Kullback-Leibler (KL) divergence, which provide direct access to the free energy difference as a variational upper or lower bound. This variational training scheme has been successfully applied to a broad range of physical problems, including lattice models Li and Wang (2018); Wu et al. (2019), atomic solids Wirnsberger et al. (2022); Ahmad and Cai (2022), quantum dots Xie et al. (2022), uniform electron gases Xie et al. (2023b), and also dense hydrogen Xie et al. (2023a). However, it still has some drawbacks for the present purpose. First, the training process typically requires a large number of energy calculations of the underlying system, which would be time-consuming for ab initio methods like DFT. As a result, many works have used empirical Wirnsberger et al. (2020, 2022); Ahmad and Cai (2022) or machine-learned potentials Molina-Taborda et al. (2024), which would inevitably introduce new source of errors. Second, symmetry constraints often play an essential role in the efficient training of flow model, especially for homogeneous systems like pure hydrogen. Although several neural network designs have been proposed to meet various symmetry requirements Wirnsberger et al. (2020, 2022), they are still not satisfactory enough regarding a proper balance between the expressive power and computational cost. The continuous normalizing flow (CNF) Chen et al. (2018); Zhang et al. (2018) is arguably the most convenient and flexible candidate now to incorporate symmetries, but its variational training requires expensive numerical integrations of ordinary differential equation (ODE) and scales unfavorably for large system sizes.
The recently developed flow matching method Albergo and Vanden-Eijnden (2023); Lipman et al. (2023); Liu et al. (2023) has largely overcome these difficulties. More specifically, one can smoothly connect the two end probability distributions by introducing an interpolant , where is a continuous “time” parameter, such as
@@
Notice this expression requires only samples from the two end distributions; there are not any on-the-fly evaluations of configurational energy or ODE integration during the training stage. On the other hand, these types of costly computations are indeed present when estimating the free energy difference after the training is finished, as shown in Eqs. (S4) and (S5). In particular, the transformed coordinates and corresponding log-Jacobian determinant should be evaluated by jointly integrating the following set of ODEs Chen et al. (2018); Zhang et al. (2018):
@@
As mentioned above, compared to previous network designs for discrete flow model Wirnsberger et al. (2020, 2022), the incorporation of various symmetries can also be more convenient for continuous normalizing flow. Specifically, the invariance of probability distribution can simply be achieved by requiring the underlying velocity field to be equivariant under the desired symmetry operation :
To do this in practice, we can leverage many recent advances in various areas, such as natural language processing Vaswani et al. (2017), molecular simulation Köhler et al. (2019); Wirnsberger et al. (2020), graph neural networks Satorras et al. (2021), and quantum many body computations Taddei et al. (2015); Pfau et al. (2020); Xie et al. (2023b); Pescia et al. (2022); Xie et al. (2023a); Wilson et al. (2023); Cassella et al. (2023); Pescia et al. (2024).
@@ Appendix C Implementation details
The MD generation of coordinate samples used for optimizing Eq. (S7) and the energy computation appearing in the free energy estimators Eq. (S4) are both performed in the DFT level using the QUANTUM ESPRESSO code Giannozzi et al. (2009). We use the exchange-correlation functional of Perdew, Burke and Ernzerhof (PBE) Perdew et al. (1996) and the pseudopotential of projector-augmented wave Blöchl (1994) type. We choose the system size , and the -space integration is performed on a Monkhorst-Pack grid Monkhorst and Pack (1976). The energy cutoff for the plane wave basis set and charge densities are and Ry, respectively. Note the parameters above have been checked to yield well converged pressures and energies. For MD simulations, the ion temperature is controlled using stochastic-velocity rescaling Bussi et al. (2007) and the time step is set to 12 a.u..
The network we use for the velocity field is adapted from FermiNet Pfau et al. (2020), which is permutation and translation equivariant by construction. We also modify the pair distance features to comply with the periodic nature of the simulation box Xie et al. (2023b); Pescia et al. (2022); Wilson et al. (2023); Cassella et al. (2023). We choose a simple linear interpolating function
with the caveat in mind that we are working on a periodic box. This setting turns out to be equivalent to the Riemannian flow matching on a torus as proposed in Ref. Chen and Lipman (2024). To allow for more efficient use of the training data, we choose to further incorporate permutation symmetry explicitly into the flow matching objective Eq. (S7) itself Klein et al. (2023). In practice, this can be achieved by finding the permutation with minimal distance for each pair of samples using the Hungarian algorithm Kuhn (1955).
@@ Appendix D Illustrative examples of the results
After the training is complete, one can further use Eq. (S4) to estimate the upper and lower bound of free energy difference between the two states, as shown by the red and black vertical line in Fig. S2, respectively. The distance between these two bounds is roughly within their own error bars, which implies our estimation have been tight enough. Figure S2 also shows histograms for the distribution of forward and reverse work , as defined in Eq. (S5). The significant overlap between these two work distributions indicates the high quality of the trained velocity field in effectively connecting the two end states being considered. Furthermore, one can clearly see the estimated free energy difference values coincide well with the intersection point of the two work distributions, where the probability densities are equal. This is a direct consequence of the fluctuation theorem as derived in Ref. Hahn and Then (2009).
We also use the same workflow outlined above to calculate the absolute free energy and entropy of the reference state at and , as mentioned in the main text. In practice, we choose the “tractable state” to correspond to a simple uniform distribution. The resulting radial distribution functions and free energy bounds are shown in Figures S3 and S4, respectively. Notice that in the present case, the two end states differ more substantially than those in Fig. S1, making the flow matching training of the velocity field more challenging. As a result, the forward and reverse work distributions also have less overlap, and the upper and lower bounds for the free energy difference are less tight compared to Fig. S2. Nonetheless, the absolute entropy of the reference state can still be estimated with a relative error of , which is comparable to the results reported by Morales et al. Morales et al. (2010a) using Hamiltonian thermodynamic integration and already satisfactory for the present purpose (see Fig. 1(a) in the main text).